# Nonlinear Causality in Brain Networks: With Application to Motor Imagery vs Execution

### v1

@@ 3.4 Connectivity at Motor Movement
    Both Fists
    [Uncaptioned image] [Uncaptioned image]
    Both Feet
    [Uncaptioned image] [Uncaptioned image]
    Table 5: Connectivity Nets at Motor Movement
    MM right-hand fist gesture (MM-RH): The MM-RH gesture exhibits more balanced connectivity between the channels. The TCI values (Table 5) show significant connectivity from AF3 and AF4 to both C3 and C4, with TCI values ranging from 7.14 to 14.29 (red arrows). The reverse connections, particularly from C3 to AF3 (TCI = 12.86, blue arrow) and from C3 to AF4 (TCI = 12.86, blue arrow), indicate a bidirectional connectivity.
    TGCI results for MM-RH reveal strong causal connectivity from AF3 to C4, with TGCI values of 18.33 (red arrow). This suggests that during right-hand fist movement, channels AF3 and AF4 play a role in driving motor activity. However, the reverse causal influence from C3 and C4 to Af3 and AF4 is less pronounced, as indicated by the lower TGCI values, such as 5.95 (blue arrow) from C3 to AF3. This pattern highlights the prominence of the prefrontal cortex in controlling motor functions during right-hand fist movement, reflecting the cognitive demand associated with executing movements of the dominant hand.
    MM both hand’s fists gesture (MM-BHs): For both fists’ gesture, the TCI results (Table 5) show relatively proportional connectivity across the prefrontal and motor cortices. The TCI values from AF4 to C3 is 14.29 (red arrow), and from AF3 to C4, it is 13.33 (red arrow), suggesting substantial connectivity over C3 and C4.
    However, the TGCI results for the both fists’ gesture seem to oppose this finding, with significant causal influence observed from C4 to AF4 (TGCI = 12.14, blue arrow) and from C3 to AF3 (TGCI = 11.90, blue arrow). The contrast in terms of TCI and TGCI during both fists’ movement suggests there is a continuous exchange of information between the channels over prefrontal and motor regions, which might be related to coordinating bilateral motor activity. The bidirectional causal connectivity suggests that both hemispheres are actively engaged, with the prefrontal cortex contributing to motor planning and the motor cortex providing feedback to adjust and refine the movements.
    MM both feet gesture (MM-BF): The MM both feet’ gesture shows the highest TCI values (Table 5) among all gestures, particularly for the connections from C3 to AF3, with a TCI of 19.29 (blue arrow), and from AF3 to C4, with a TCI of 14.29 (red arrow). These results suggest that the MM both feet’ task involves strong connectivity between the prefrontal and motor cortices, especially between the left prefrontal and both motor areas.
@@
    The TGCI results for MM-BF support this interpretation, showing very high TGCI values, especially from C3 to AF3 (TGCI = 31.67, blue arrow). This indicates a principal causal influence of the C3 over the AF3 during MM-BF, highlighting the significant feedback mechanisms necessary for executing and adjusting bilateral feet movement. The causal influence from AF3 to C3 and C4 is also strong, particularly from AF3 to C4 (TGCI = 19.29, red arrow), reflecting the prefrontal cortex’s role in controlling these complex movements.
    Comparative Analysis Between gestures: Across all gestures, several key trends emerge. First, the overall connectivity strength, as measured by TCI, tends to be higher for both fists’ and both feet’ gestures than the left-hand fist and right-hand fist gestures. This suggests that bilateral motor tasks require more extensive neural coordination, with stronger interactions between the prefrontal and motor cortices.
    Second, the TGCI results indicate that during unilateral tasks (left-hand fist and right-hand fist), the motor cortex wields a more prominent causal influence over the prefrontal cortex, especially in the left-hand fist gesture. This pattern likely reflects the need for real-time feedback and adjustments during the execution of these movements, particularly when performing tasks with the non-dominant hand (left-hand fist).
    Finally, the bidirectional connectivity observed in TGCI results during both fists’ and both feet’ gestures underlines the importance of continuous feedback and coordination between the channels over prefrontal and motor regions during these complex bilateral tasks. This feedback mechanism suggests that both hemispheres are involved in synchronizing movements, which appears to be necessary for the successful execution of coordinated bilateral actions.
    The connectivity patterns across the different gestures highlight the varying demands on neural coordination and causal influence depending on the specific motor task. Bilateral movements (both fists’ and both feet’) involve stronger and more complex connectivity, reflecting the higher cognitive and motor demands associated with coordinating movements across both sides of the body. Unilateral movements, particularly those involving the non-dominant hand, require significant motor-to-prefrontal feedback, emphasizing the importance of real-time adjustments during these tasks.
    3.5 Comparative Analysis of Threshold Granger Causality: Motor Imagery vs Motor Movement
@@
    We have thus far presented and discussed the connectivity between the channels separately for MI and MM tasks using TCI and TGCI values. The figures presented in Table 4 and Table 5 visualize the corresponding results. Furthermore, we conducted an additional analysis to assess the differences in causal connectivities within and between MI and MM tasks, employing a permutational Hotelling’s $T^{2}$ test (Curran and Hersh (2021)) on data from 35 subjects. In these tests, the connectivity of each pair within MI and MM was compared with the connectivity of the same pair in the opposite direction. To illustrate, the TGC sampling distributions of MI/MM at $X$ to $Y$ were tested against the TGC sampling distributions of MI/MM at $Y$ to $X$. The figures given in Table 7 display the significant differences in connectivities during MI and those during MM in columns 1 and 2, respectively. The differences between MI and MM were estimated by comparing the connectivity distributions in the same direction and provided by column 3. That is, the TGC sampling distributions at $X$ to $Y$ during MI were compared to the TGC sampling distributions during MM at $X$ to $Y$. Each arrow in the table signifies a statistically significant difference in directional connectivity, with the color indicating the direction of influence: red arrows represent prefrontal-to-motor differences, blue arrows indicate motor-to-prefrontal differences, green arrows signify stronger connectivity in favor of MI, purple arrows indicate stronger connectivity in favor of MM. The implemented procedure followed during the comparison is outlined in Table 6:
@@
    For all subject; Let $X$ and $Y$ represent two nodes in the EEG network. 1. Apply the threshold nonlinearity test using the 12th lag TAR model for each epoch and combine test results. Then, generate a 35 (i.e., number of subjects) by 1 vector for each delay $d$. (a) Retain the F-test statistics for subjects where linearity is rejected and assign 0s for others. 2. Apply the Wald test on the data of subjects who reject linearity to assess the predictive power of the threshold variable $X_{t-d}$. (a) Filter the vector of F-test statistics based on subjects who pass the Wald test. (b) Retain only those elements corresponding to subjects who pass the Wald test, forming the TGC samples for each delay $d$. 3. Compare the TGC sampling distributions for each pairwise channel interaction $X$ to $Y$ vs $Y$ to $X$ within the same task (i.e., MI or MM) using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in the directional connectivity between $X$ to $Y$ vs $Y$ to $X$ within the same task (i.e., MI or MM). (b) Visualize significant differences with arrows: Red arrows for significant prefrontal-to-motor connectivity. Blue arrows for significant motor-to-prefrontal connectivity. 4. Compare the TGC sampling distributions in the same direction (e.g., $X$ to $Y$) across MI and MM tasks using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in connectivity between MI and MM for each channel pair. (b) Visualize significant differences with arrows: Green arrows for stronger connectivity at MI compared to MM. Purple arrows for stronger connectivity at MM compared to MI.
    Table 7: Threshold Granger Causality Differences for Motor Imagery (MI), Motor Execution (MM), and MI vs MM across different motor tasks
    Differences at MI Differences at MM Differences MIvsMM
    Left Fist
@@
    Both Fists
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    Both Feet
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    For the left-hand fist MI task, the statistically significant differences in TGC samples given in Table 7 show bidirectional connectivities between AF4 and C4, with a red arrow from AF4 to C4 and a blue arrow from C4 to AF4. The TGCI results support this finding, indicating a strong connection from AF4 to C4 (TGCI = 25.24, red arrow) and notable feedback from C4 to AF4 (TGCI = 13.10, blue arrow). This consistency between the significant differences in TGC samples and TGCI results highlights bidirectional connectivity between these channels during MI, emphasizing the role of AF4 in influencing the C4 and the feedback mechanism. Table 7 also shows a significant connection from C4 to AF3 (blue arrow) for the MI-LH, which is indeed captured by the TGCI results, with a TGCI value of 12.62 for C4 to AF3 and 5.95 for AF3 to C4. This demonstrates that the connection from C4 to AF3 is more prominent than the reverse, indicating that C4 exerts more influence on the prefrontal region AF3 during MI-LH.
    In the MM task for left-hand fist movement, both the significant differences in TGC samples and TGCI results emphasize strong motor-to-prefrontal connectivity, particularly from C3 and C4 to AF3 (blue arrows). The TGCI values are particularly high from C3 to AF3 (TGCI = 25.00, blue arrow), highlighting the motor area channels’ significant influence on AF3 during actual movement. The red arrows from AF3 and AF4 to C4 in Table 7 may indicate the low degree of influence of the prefrontal area channels over C4 since it is not highlighted by TGCI.
    For the right-hand fist MI task, Table 7 shows significant motor-to-prefrontal influence, with blue arrows from C3 to AF3 and AF4. This dynamic does not appear to align with the TGCI outcomes, wherein C4 exerts an influence on AF3 and AF4 (TGCI = 6.19, indicated by the blue arrow). Additionally, TGCI results highlight strong prefrontal-to-motor connectivity, particularly from AF3 and AF4 to C4 (TGCI = 18.33 for both, red arrows), suggesting that during MI, AF3, and AF4 play a notable role in driving motor area channels with considerable feedback from C3 and C4. The discordance between the statistically significant differences in TGC samples and TGCI outcomes suggests that both methodologies provide complementary insights into the bidirectional connectivity essential for the intricate cognitive process of motor imagery.
    In the right-hand fist MM task, Table 7 shows significant connectivity from AF4 to C4 (red arrow), which corresponds with the TGCI results, where AF4 to C4 is also highlighted (TGCI = 6.43, red arrow). This alignment between the significant differences in TGC samples and TGCI suggests that AF4 may play a critical role in driving RH motor activity. However, the feedback from C3 to AF4 identified in Table 7 (blue arrow) does not have a direct counterpart in the TGCI results. Additionally, the TGCI results indicate a strong connection from AF3 to C4 (TGCI = 18.33, red arrow) and also a notable connection from C4 to AF3 (TGCI = 11.9, blue arrow). This finding suggests that there is significant feedback from C4 to AF3, which may play a critical role in RH motor activity. This difference indicates that while TGCI captures the influence from the prefrontal area channels to motor area channels, the significant differences in TGC samples highlight a refined bidirectional interaction, particularly from channels over the motor area to the prefrontal area, that may not be as prominent in the TGCI analysis.
@@
    For both fists’ MI task, Table 7 highlights significant connectivity from AF3 and AF4 to C4 (red arrows), which is consistent with the TGCI results showing similar connections (TGCI = 12.14 for AF3 to C4, red arrow). This alignment between the significant differences in TGC samples and TGCI underlines the importance of AF3 and AF4 in coordinating both fists’ movement during MI. Table 7 also identifies significant motor-to-prefrontal feedback from C3 to AF4 (blue arrow), and this is confirmed by TGCI results, which show a strong connection from C3 to AF4 (TGCI = 17.86, blue arrow). The agreement between the two methods suggests robust bidirectional dynamics during both fists’ MI, with information flow between channels over the motor and prefrontal parts.
@@
    In both fists’ MM task, Table 7 identifies connectivity from AF3 to C3 and AF4 to C3 (red arrows). This finding is consistent with the TGCI results, which also highlight connections from AF3 to C4 (TGCI = 6.19, red arrow) and AF4 to C4 (TGCI = 5.95, red arrow). These connections emphasize the critical role of AF3 and AF4 in coordinating both fists’ movement during MM. However, the blue arrows from C3 to AF3 and C3 to AF4 in Table 7 indicate significant feedback from C3 to AF3 and AF4, which are partially captured by TGCI results, such as the connection from C4 to AF3 (TGCI = 11.9, blue arrow). This feedback highlights the bidirectional nature of the connectivity necessary for maintaining coordination during bilateral fists movement. The combination of TGC sample differences and TGCI results highlights the condition for continuous, reciprocal communication between the prefrontal and motor regions to achieve coordinated motor control.
@@
    For both feet’ MI task, Table 7 shows significant differences from C3 to AF3 and C3 to AF4 (blue arrows). These connections indicate that C3 influences both the left and right prefrontal area channels during this task. Additionally, Table 7 reveals a significant connection between AF3 and C4 (red arrow), indicating that the AF3 influence over the C4 was evident during imagery of bilateral foot movements. Comparing these findings with the TGCI results, we observe a strong connection from C3 to AF3 (TGCI = 12.14, blue arrow) and from C3 to AF4 (TGCI = 24.05, blue arrow), reinforcing the importance of these pathways during MI. There is also a consistent outcome with regard to the connectivity between AF3 and C4 (TGCI = 6.19, as indicated by the red arrow in Table 7). The combination of both emphasizes the bidirectional dynamics necessary for coordinating complex motor imagery tasks involving both feet’.
@@
    In the MM task for both feet’, Table 7 shows significant connectivity from AF3 to C3, AF3 to C4, and AF4 to C4 (red arrows), highlighting the prominent role of AF3 and AF4 in driving motor activity during actual bilateral foot movement. Additionally, significant feedback is observed from C3 to AF3 (blue arrow). TGCI results for both feet’ MM task further emphasize this feedback, particularly highlighting the strong connection from C3 to AF3 (TGCI = 31.67, blue arrow), which is the most pronounced feedback pathway observed. The consistency between the significant differences in TGC samples and TGCI findings suggests that both methods effectively capture the critical bidirectional flow of information between the channels over the prefrontal and motor areas during MM. The significant connections identified by TGCI (e.g., AF3 to C4 with TGCI = 19.29, red arrow and AF4 to C3 with TGCI = 11.9, red arrow) further support the argument that the AF3 and AF4 take a part in the bilateral motor activity required for both feet’ movement.
    MI vs. MM Task Comparisons: In the left-hand fist movement, Table 7 reveals an interplay between MI and MM. The green arrows from AF4 to C4 and C4 to AF4 suggest that MI engages feedback connectivity between the right prefrontal and right motor compared to MM, which is consistent with the TGCI results showing significant connectivity from AF4 to C4 (TGCI = 25.24, red arrow) and C4 to AF4 (TGCI = 13.1, red arrow) during MI. This indicates that AF4 and C4 are more strongly engaged during MI for left-hand fist movement than during MM. On the other hand, the purple arrows from AF3 to C3, from AF3 to C4, from C3 to AF3, from C3 to AF4, and from C4 to AF3 in Table 7 suggest that MM necessitates enhanced connectivity between these channels when compared to MI. This pattern aligns with the TGCI findings that emphasize the importance of feedback during actual movement, suggesting that during MM, C3, and C4 have a more prominent influence on AF3 and AF4.
    For the right-hand fist task, the significant differences in TGC samples in Table 7 show a mixed pattern of connectivity between MI and MM. The green arrows from C3 to AF4 and from C4 to AF4 suggest that MI relies more heavily on motor-to-prefrontal influence. This observation contrasts with the TGCI results, where MI shows significant prefrontal-to-motor connectivity, particularly from AF3 and AF4 to C4 (TGCI = 18.33, red arrow). Similar discordance was observed between the statistically significant differences in TGC samples and TGCI outcomes in the right-hand fist MI task. On the other hand, the purple arrows from AF3 to C4, AF4 to C3, and C4 to AF3 in Table 7 indicate that MM involves stronger prefrontal-to-motor connectivity, slightly aligns with the TGCI results that highlight AF3 to C4 (TGCI = 18.33, red arrow) and AF4 to C4 (TGCI = 6.43, red arrow) as critical connections during right-hand fist movement.
@@
    In both fists’ movement, Table 7 column 3 shows a clear preference for MI in terms of bidirectional connectivity, as indicated by the green arrows from AF3 to C4, AF4 to C3, C3 to AF4, C4 to AF3, and C4 to AF4. These findings suggest that MI requires a more integrated and reciprocal interaction between channels over prefrontal and motor regions to coordinate bilateral fist movements. The TGCI results support this observation, with significant connectivity noted from C3 to AF4 (TGCI = 17.86, blue arrow) and from C4 to AF3 (TGCI = 11.90, blue arrow) during MI. Conversely, the absence of purple arrows in Table 7 for both fists’ movement indicates that MM does not engage these bidirectional connections as strongly as MI, suggesting a more straightforward connectivity pattern during actual movement.
@@
    In both feet’ movement task, Table 7 highlights a strong preference for MI, as evidenced by the green arrows indicating stronger bidirectional connectivity from AF4 to C3, AF4 to C4, C3 to AF3, C3 to AF4, C4 to AF3, and C4 to AF4. These results suggest that MI of both feet’ involves complex interactions between the channels over prefrontal and motor areas, likely due to the cognitive demands of coordinating bilateral foot movements. The TGCI results corroborate this, with strong feedback from C3 to AF4 (TGCI = 24.05, blue arrow) and from C4 to AF4 (TGCI = 18.10, blue arrow) during MI. The purple arrow from AF3 to C4 in Table 7 suggests that MM involves stronger prefrontal-to-motor connectivity in this specific pathway, aligning with the TGCI result showing significant connectivity from AF3 to C4 (TGCI = 19.29, red arrow) during MM. This difference in connectivity patterns between MI and MM for both feet’ motion highlights the distinct neural circuitries employed for imagined versus actual movement, with MI requiring more extensive feedback and coordination across hemispheres.
@@
    In summary, the comparison between MI and MM across different tasks reveals that MI generally engages in more complex and bidirectional connectivity, particularly in tasks involving bilateral coordination. This is reflected in the significant differences in TGC samples and supported by TGCI results, which together provide an important view of the distinct neural mechanisms underlying motor imagery and motor movement. The significant differences in TGC samples are particularly sensitive to bidirectional feedback dynamics, capturing refined interactions that are also reflected in TGCI results, though sometimes with different emphases. TGCI, on the other hand, provides a clear quantification of the overall strength and directionality of connectivity, highlighting key connections that drive motor control and imagery. In light of the two distinct perspectives from which these connections can be evaluated, it seems more analytical to perceive seemingly unrelated dissimilarities as complementary rather than exclusive.

### v2

@@ 3.4 Connectivity at Motor Execution
    Table 5: Connectivity Nets at Motor Execution
@@
    Motor execution - both feet gesture (ME-BF). The ME-BF gesture shows the highest TCI values (Table 5) among all gestures, particularly for C3 $\to$ AF3 19.29 (blue arrow), and AF3 $\to$ C4 14.29 (red arrow). Findings suggest that ME-BF involves strong connectivity between the prefrontal and motor cortices, especially between the left prefrontal and both motor areas. The TGCI results support this interpretation, showing very high TGCI values, especially C3 $\to$ AF3 31.67 (blue arrow). This indicates a principal causal influence of the C3 over the AF3 during ME-BF, highlighting the significant feedback mechanisms necessary for executing and adjusting bilateral feet execution. A strong influence was observed AF3 $\to$ C4 19.29 (red arrow), reflecting the prefrontal cortex’s role in the execution of this bilateral movement.
    Comparative analysis between gestures. The ME-BF gesture shows stronger connectivity than ME-BHs, particularly in motor-to-prefrontal feedback. Higher TCI values (C3 $\to$ AF3: 19.29 vs. 5.95; C4 $\to$ AF3: 13.33 vs. 6.19) indicate greater motor cortex control over prefrontal areas during ME-BF, especially in the left hemisphere. In contrast, ME-BHs show weaker motor-to-prefrontal feedback, with more prefrontal control over the motor cortex. Both ME-BHs and ME-BF gestures involve prefrontal control over motor activity, but motor-to-prefrontal feedback is stronger in ME-BF, particularly from C3 $\to$ AF3. Cross-hemispheric connectivity is significant in both gestures, but stronger in ME-BF, particularly between C3, AF4, C4, and AF3. This reflects the greater bilateral coordination needed for ME-BF compared to ME-BHs movements. Both gestures show bidirectional connectivity between prefrontal and motor cortex, with stronger motor-to-prefrontal feedback in ME-BF. The important difference is the dominant role of the motor cortex in ME-BF, especially in motor-to-prefrontal regulation, and stronger cross-hemispheric interactions during ME-BF.
    3.5 Comparative analysis of threshold Granger causality: imagery vs. execution
@@
    We have thus far discussed the connectivity for MI and ME using TCI and TGCI values, as visualized in Tables 4 and 5. Further, we performed a permutational Hotelling’s $T^{2}$ test (Curran and Hersh (2021)) on data from 35 subjects to examine causal connectivity differences within and between MI and ME tasks. For each pair, connectivity in one direction (e.g., $X$to $Y$) was compared to the reverse direction ($Y$to $X$) within MI and ME tasks. Significant differences are shown in Table 7, with columns 1 and 2 highlighting MI and ME differences, and column 3 showing differences between MI and ME in the same direction ($X$to $Y$). Each arrow represents a statistically significant directional difference: red for prefrontal-to-motor, blue for motor-to-prefrontal, green for MI prominence, and purple for ME prominence. The comparison procedure is outlined in Table 6:
@@
    For all subject; Let $X$ and $Y$ represent two nodes in the EEG network. 1. Apply the threshold nonlinearity test using the TAR(12) model for each epoch and combine test results. Then, generate a 35 (i.e., number of subjects) by 1 vector for each delay $d$. (a) Retain the F-test statistics for subjects where linearity is rejected and assign 0s for others. 2. Apply the Wald test on the data of subjects who reject linearity to assess the predictive power of the threshold variable $X_{t-d}$. (a) Filter the vector of F-test statistics based on subjects who pass the Wald test. (b) Retain only those elements corresponding to subjects who pass the Wald test, forming the TGC distribution for each delay $d$. 3. Compare the TGC distributions for each pairwise channel interaction $X$to $Y$ vs $Y$to $X$ within the same task (i.e., MI or ME) using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in the directional connectivity between $X$to $Y$ vs $Y$to $X$ within the same task (i.e., MI or ME). (b) Visualize significant differences with arrows: Red arrows for significant prefrontal-to-motor connectivity. Blue arrows for significant motor-to-prefrontal connectivity. 4. Compare the TGC distributions in the same direction (e.g., $X$to $Y$) across MI and ME tasks using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in connectivity between MI and ME for each channel pair. (b) Visualize significant differences with arrows: Green arrows for stronger connectivity at MI compared to ME. Purple arrows for stronger connectivity at ME compared to MI.
    Differences at MI Differences at ME Differences MIvsME
    Both Fists (BHs)
@@
    Table 7: Threshold Granger Causality Differences for Motor Imagery (MI), Motor Execution (ME), and MI vs ME across different motor tasks
@@
    For the MI-BHs, Table 7 highlights significant connectivity from AF3 $\to$ C4 and AF4 $\to$ C4 (red arrows), which is consistent with the TGCI results showing similar connections (TGCI = 12.14 for AF3 $\to$ C4, red arrow). This alignment between the significant differences in TGC samples and TGCI underlines the importance of AF3 and AF4 in coordinating BHs movement during MI. Table 7 also identifies significant motor-to-prefrontal feedback for MI-BHs from C3 $\to$ AF4 (blue arrow), and this is confirmed by TGCI results, which show a strong connection from C3 $\to$ AF4 (TGCI = 17.86, blue arrow). The agreement between the two methods suggests robust bidirectional dynamics during MI-BHs, with information flow between channels over the motor and prefrontal parts.
@@
    In the ME-BHs, Table 7 identifies connectivity from AF3 $\to$ C3 and AF4 $\to$ C3 (red arrows). This finding is consistent with the TGCI results, which also highlight connections from AF3 $\to$ C4 (TGCI = 6.19, red arrow) and AF4 $\to$ C4 (TGCI = 5.95, red arrow). These connections emphasize the critical role of AF3 and AF4 in coordinating the ME-BHs. However, the blue arrows from C3 $\to$ AF3 and C3 $\to$ AF4 in Table 7 indicate significant feedback from C3 $\to$ AF3 and $\to$ AF4, which are partially captured by TGCI results, such as the connection from C4 $\to$ AF3 (TGCI = 11.9, blue arrow). This feedback highlights the bidirectional nature of the connectivity necessary for maintaining coordination during bilateral fists task. The combination of TGC sample differences and TGCI results highlights the condition for continuous, reciprocal communication between the prefrontal and motor regions to achieve coordinated motor execution.
@@
    For the MI-BF, Table 7 shows significant differences from C3 $\to$ AF3 and C3 $\to$ AF4 (blue arrows). These connections indicate that C3 influences both the left and right prefrontal area channels during this task. Additionally, Table 7 reveals a significant connection between AF3 and C4 (red arrow), indicating that the AF3 influence over the C4 was evident during bilateral MI-BF. Comparing these findings with the TGCI results, we observe a strong connection from C3 $\to$ AF3 (TGCI = 12.14, blue arrow) and from C3 $\to$ AF4 (TGCI = 24.05, blue arrow), reinforcing the importance of these pathways during MI. There is also a consistent outcome with regard to the connectivity between AF3 and C4 (TGCI = 6.19, as indicated by the red arrow in Table 7). The combination of both emphasizes the bidirectional dynamics necessary for coordinating complex motor imagery tasks involving both feet.
@@
    In contrast, for the ME-BF, Table 7 shows significant connectivity from AF3 $\to$ C3, AF3 $\to$ C4, and AF4 $\to$ C4 (red arrows), highlighting the prominent role of AF3 and AF4 in driving motor activity during bilateral ME-BF. Additionally, significant feedback is observed from C3 $\to$ AF3 (blue arrow). TGCI results for the ME-BF further emphasize this feedback, particularly highlighting the strong connection from C3 $\to$ AF3 (TGCI = 31.67, blue arrow), which is the most pronounced feedback pathway observed. The consistency between the significant differences in TGC samples and TGCI findings suggests that both methods effectively capture the critical bidirectional flow of information between the channels over the prefrontal and motor areas during ME. The significant connections identified by TGCI (e.g., AF3 $\to$ C4 with TGCI = 19.29, red arrow and AF4 $\to$ C3 with TGCI = 11.9, red arrow) further support the argument that the AF3 and AF4 take a part in the bilateral ME-BF.
    Comparison: imagery vs. execution
@@
    Table 7, 1st-row 3rd-column show a clear preference in favor of MI-BHs regarding bidirectional connectivity, as indicated by the green arrows from AF3 $\to$ C4, AF4 $\to$ C3, C3 $\to$ AF4, C4 $\to$ AF3, and C4 $\to$ AF4. These findings suggest that MI requires a more integrated and reciprocal interaction between channels over prefrontal and motor regions to coordinate bilateral fists movements. The TGCI results support this observation, with significant connectivity noted from C3 $\to$ AF4 (TGCI = 17.86, blue arrow) and from C4 $\to$ AF3 (TGCI = 11.90, blue arrow) during MI. Conversely, the absence of purple arrows in Table 7 for both fists movement indicates that ME does not engage these bidirectional connections as strongly as MI, suggesting a more straightforward connectivity pattern during actual movement.
@@
    Table 7, 2nd-row 3rd-column highlights a strong preference for MI-BF, as evidenced by the green arrows indicating stronger bidirectional connectivity from AF4 $\to$ C3, AF4 $\to$ C4, C3 $\to$ AF3, C3 $\to$ AF4, C4 $\to$ AF3, and C4 $\to$ AF4. These results suggest that MI of BF involves complex interactions between the channels, likely due to the cognitive demands of coordinating bilateral feet movement. The TGCI results corroborate this, with strong feedback from C3 $\to$ AF4 (TGCI = 24.05, blue arrow) and from C4 $\to$ AF4 (TGCI = 18.10, blue arrow) during MI. The purple arrow from AF3 $\to$ C4 in Table 7, 2nd-row 3rd-column suggests that ME involves stronger prefrontal-to-motor connectivity in this specific pathway, aligning with the TGCI result showing significant connectivity from AF3 $\to$ C4 (TGCI = 19.29, red arrow) during ME. This difference in connectivity patterns between MI and ME for BF motion tasks highlights the distinct neural circuitries employed for imagined versus execution, with MI requiring more extensive feedback and coordination across hemispheres.
@@
    In summary, the comparison between MI and ME across different tasks reveals that MI generally engages in more complex and bidirectional connectivity, particularly in tasks involving bilateral coordination. This is reflected in the significant differences in TGC samples and supported by TGCI results, which together provide an important view of the distinct neural mechanisms underlying MI and ME. The significant differences in TGC samples are particularly sensitive to bidirectional feedback dynamics, capturing refined interactions that are also reflected in TGCI results, though sometimes with different emphases. TGCI, on the other hand, provides a clear quantification of the overall strength and directionality of connectivity, highlighting key connections that drive motor control and imagery. In light of the two distinct perspectives from which these connections can be evaluated, it seems more analytical to perceive seemingly unrelated dissimilarities as complementary rather than exclusive.